OpenStudy (anonymous):

Find the Maximum y-value on the f(x) f(x)= -x^2+4x+9

OpenStudy (anonymous):

@HawkCrimson I can't remember if you helped me last time or not?

OpenStudy (mertsj):

Do you recognize that this is a parabola which is concave downward?

OpenStudy (anonymous):

No?

OpenStudy (anonymous):

leading coefficient is negative ( it is -1) so the parabola opens down 

OpenStudy (anonymous):

okay, that makes sense but putting that in a formula how do you figure it out? by x= -b/2a?

OpenStudy (anonymous):

maximum is the second coordinate of the vertex the first coordinate of the vertex of \(y=ax^2+bx+c\) is always \(-\frac{b}{2a}\)

OpenStudy (anonymous):

in your case it is \(-\frac{4}{2\times (-1)}=2\)

OpenStudy (anonymous):

second coordinate of the vertex, which is what you are looking for, is what you get when you replace \(x\) by \(2\)

OpenStudy (anonymous):

-4/ 2(4)= -4/6 = -2/3? Or is that completely wrong?

OpenStudy (anonymous):

or do you bring it back to the equation of 2^2+4*(2)+9

OpenStudy (anonymous):

which if I am correct gives you a max of 25

OpenStudy (mertsj):